A sample of nitrogen gas was collected over water at 30oC at which temperature the saturated vapour pressure of water is 4.26kPa. 150mL of gas was collected at a total pressure of 66.7kPa. Calculate the number of moles of dry nitrogen collected.
1
Expert's answer
2020-09-21T06:30:37-0400

A sample of nitrogen gas was collected over water at 30oC at which temperature the saturated vapour pressure of water is 4.26kPa. 150mL of gas was collected at a total pressure of 66.7kPa.


So , Total Pressure = 66.7 kPa .

Vapour pressure of water = 4.2 kPa .


Pressure of dry gas ( P ) = (66.7 - 4.2) kPa = 62.5 kPa . = ( 62.5 * 1000 ) pa

= ( 62500 / 100000 ) atm . = 0.625 atm .



Volume ( V ) = 150 ml = ( 150 /1000 ) liter = 0.150 liter .


Temperature ( T )= 30 degree = ( 273 +30 ) k = 303 k .


P V = n R T.


( 0.625 atm) * ( 0.150 liter ) = n * ( 0.0821 atm . liter / mol . k ) * 303 k


So , n = 0.00376 mole . = moles of dry nitrogen gas .